# Multisoliton solutions for equivariant wave maps on a $2+1$ dimensional wormhole

### v1

@@ 1. Introduction
While most of the mathematical work on soliton resolution conjecture has been concerned with nonlinear radial wave equations on $d+1$ Minkowski spacetime, it was pointed out by the first author [5] that the problem becomes more tractable for the corresponding equations posed on a wormhole spacetime. Subsequently, the soliton resolution conjecture for equivariant wave maps from the $3+1$ dimensional wormhole into the 3-sphere was formulated and verified numerically in [6] and later proved by Rodriguez [7, 8]. The simplicity of this model stems from the fact that in each topological sector there is a unique endstate of evolution given by the rigid stable soliton. In this paper we consider the analogous model in two spatial dimensions, i.e. equivariant wave maps from the $2+~{}1$ dimensional wormhole into the 2-sphere. This model admits an explicit static solution (harmonic map of degree one, hereafter called the kink), whose reflections (called antikinks) and conformal transformations serve as building blocks for a variety of multisoliton solutions. Here, as the first step towards the taxonomy of possible endstates, we consider multisoliton solutions of degree zero and one which are asymptotically static. Although such “parabolic” solutions are not generic, they play are important role as separatrices between different generic behaviours in the global phase portrait.
The paper is organized as follows. In section 2 we introduce the model and formulate our main conjecture asserting the existence of asymptotically static chains of $N$ alternating kinks and antikinks whose positions and mutual distances tend to infinity as $t\rightarrow\infty$. In section 3 we derive an effective finite dimensional system of ordinary differential equations that approximate the asymptotic dynamics of $N$-chains. Finally, in section 4 we present the results of direct numerical computations of the wave map equation and confirm the predictions of the ODE model for $N=2,3$.
@@ 2. Model
are the kinetic and potential energies. Any finite energy solution must be equal to integer multiples of $\pi$ at $x=\pm\infty$. Without loss of generality we set $u(t,-\infty)=0$, so that the integer $n=u(t,+\infty)/\pi$ determines the topological degree of solution (which is preserved in evolution). For $n=0$ the energy attains the minimum $E=0$ at the vacuum $u=0$. For $n=1$ we have the Bogomolnyi inequality
@@
$u_{N}(t,x):=$ $Q(x)+\sum\limits_{j=1}^{J}(-1)^{j}\left[Q(x+c_{j}(t))+Q(x-c_{j}(t))\right],$ $N=2J+1$ (12b)
@@ 3. ODE approximation
In this section we derive finite-dimensional ODE models for the asymptotic dynamics of the $N$-chains (12b) via the method of collective coordinates. This method works as follows; see [12] for the general exposition. The lagrangian functional $\mathcal{L}=T-V$ for the wave map equation (6) is evaluated on the $N$-chain configuration $u_{N}(t,x)$ by integrating over $x$. The resulting function is interpreted as an effective lagrangian $\mathcal{L}_{\text{eff}}$ which determines the evolution of the kink and antikink positions $c_{j}(t)$. Thus, instead of extremizing $\mathcal{L}$ with respect to all variations of $u(t,x)$, one considers only variations allowed by the ansatz. This approach omits radiative degrees of freedom so it is expected to provide good approximation to true dynamics only if radiation is negligible, which is precisely the case at hand.
@@
from which we see that the effective kinetic energy of an isolated kink positioned at some large $c(t)$ is approximately equal to $\frac{\pi}{2}e^{2c}{\dot{c}}^{2}$, while the binding energy of the kink-antikink pair separated by a large distance $d$ is approximately equal to $-16e^{-2d}$. Using this, we get the effective lagrangians which determine the asymptotic dynamics of $N$-chains (here by convention $c_{0}=0$)
@@
which is exact. For $J\geq 2$, the solution $r_{j}^{(2J+1)}(t)$ can be constructed formally as follows. Assuming temporarily that the right hand sides of equations (16) vanish for $11\leq j\leq J-$ 1, one gets $r_{j-1}r_{j+1}=r_{j}^{2}$, thus $r_{j}$ form a geometric sequence
@@
and therefore $r_{J-1}=r_{J}^{\frac{J-1}{J}}$. Substituting this into equation (16) for $j=J$ gives $\ddot{r}_{J}=-r_{J}^{-\frac{J+2}{J}},$ whose zero energy solution is $r_{J}(t)=(At)^{\frac{J}{J+1}}$. Combining this result with (20) yields (18). Having the leading order asymptotics, the solution can be systematically improved by adding subleading corrections. Let us illustrate how this goes for $J=2$, where (18) reads
@@ 4. PDE computations
In this section we solve equation (6) numerically using the hyperboloidal formulation of the initial value problem [13]. To this end we define a new time coordinate $s=t-\cosh{x}$. The hypersurfaces of constant $s$ are ‘hyperboloidal’ which means that they are spacelike hypersurfaces that approach the future null infinities of the wormhole spacetime along the outgoing null cones for $x\rightarrow\pm\infty$. In terms of $s$ equation (6) becomes (for $k=2$)
@@
where $b_{\pm}(s)$ are the radiation coefficients. Multiplying equation (36) by $u_{s}$ and rearranging terms we obtain the conservation law
@@
Note that the potential part of the energy $\mathcal{E}$ is the same as $V(u)$ for the conserved energy $E$, however the kinetic part is different. Since the Bondi energy is monotonically decreasing and bounded from below – by 00 for $n=0$ and by $4$ for $n=1$ (due to Bogomolnyi’s inequality (8)) – there exists a limit
@@
In numerical studies we introduce the coordinate $y=\tanh\left(x/4\right)$ which compactifies the spatial domain to the interval $[-1,1]$ and transforms equation (36) to
$u_{ss}+\frac{2y(1+y^{2})}{1-y^{2}}\,u_{sy}+\frac{y^{4}+6y^{2}+1}{(1-y^{2})^{2}}u_{s}=\frac{1}{16}(1-y^{2})\partial_{y}\left(1-y^{2})u_{y}\right)-2\sin(2u).$ (42)
@@
Acknowledgement. This work was supported by the National Science Centre grant #2017/26/A/ST2/00530 (PB), the ERC Starting Grant ”INSOLIT” 101117126 (JJ), the Austrian Science Fund (FWF) Project P 36455 and START-Project Y 963 (MM).

### v2

@@ 1. Introduction
While most of the mathematical work on soliton resolution conjecture has been concerned with nonlinear radial wave equations on $d+1$ Minkowski spacetime, it was pointed out by the first author [5] that the problem becomes more tractable for the corresponding equations posed on a wormhole spacetime. Subsequently, the soliton resolution conjecture for equivariant wave maps from the $3+1$ dimensional wormhole into the 3-sphere was formulated and verified numerically in [6] and later proved by Rodriguez [7, 8]. The simplicity of this model stems from the fact that in each topological sector there is a unique endstate of evolution given by the rigid stable soliton. In this paper we consider the analogous model in two spatial dimensions, i.e. equivariant wave maps from the $2+~{}1$ dimensional wormhole into the 2-sphere. This model admits an explicit static solution (harmonic map of degree one, hereafter called the kink), whose reflections (called antikinks) and conformal transformations serve as building blocks for a variety of multisoliton solutions. Here, as the first step toward the taxonomy of possible end states, we consider multisoliton solutions of degree zero and one which are asymptotically static. Although such “parabolic” solutions are not generic, they play an important role as separatrices between different generic behaviors in the global phase portrait. We note that similar solutions in the case of the $\phi^{4}$ model were considered in [9, 10].
The paper is organized as follows. In Sec. 2 we introduce the model and formulate our main conjecture asserting the existence of asymptotically static chains of $N$ alternating kinks and antikinks whose positions and mutual distances tend to infinity as $t\rightarrow\infty$. In Sec. 3 we derive an effective finite dimensional system of ordinary differential equations that approximate the asymptotic dynamics of $N$-chains. Finally, in Sec. 4 we present the results of direct numerical computations of the wave map equation and confirm the predictions of the ODE model for $N=2,3$.
@@ 2. Model
are the kinetic and potential energies. Any finite energy solution must be equal to integer multiples of $\pi$ at $x=\pm\infty$. Without loss of generality we set $u(t,-\infty)=0$, so that the integer $n=u(t,+\infty)/\pi$ determines the topological degree of solution (which is preserved in evolution). For $n=0$ the energy attains the minimum $E=0$ at the vacuum $u=0$. For $n=1$ we have the Bogomol’nyi inequality
@@
$u_{N}(t,x):=$ $Q(x)+\sum\limits_{j=1}^{J}(-1)^{j}\left[Q(x+c_{j}(t))+Q(x-c_{j}(t))\right],$ $N=2J+1$
@@ 3. ODE approximation
In this section we derive finite-dimensional ODE models for the asymptotic dynamics of the $N$-chains (2) via the method of collective coordinates. This method works as follows; see [12] for the general exposition. The Lagrangian functional $\mathcal{L}=T-V$ for the wave map equation (6) is evaluated on the $N$-chain configuration $u_{N}(t,x)$ by integrating over $x$. The resulting function is interpreted as an effective Lagrangian $\mathcal{L}_{\text{eff}}$ which determines the evolution of the kink and antikink positions $c_{j}(t)$. Thus, instead of extremizing $\mathcal{L}$ with respect to all variations of $u(t,x)$, one considers only variations allowed by the ansatz. This approach omits radiative degrees of freedom so it is expected to provide good approximation to true dynamics only if radiation is negligible, which is precisely the case at hand.
@@
from which we see that the effective kinetic energy of an isolated kink positioned at some large $c(t)$ is approximately equal to $\frac{\pi}{4}e^{2c}{\dot{c}}^{2}$, while the binding energy of the kink-antikink pair separated by a large distance $d$ is approximately equal to $-16e^{-2d}$. Using this, we get the effective Lagrangians which determine the asymptotic dynamics of $N$-chains (here by convention $c_{0}=0$)
@@
which is exact. For $J\geq 2$, the solution $r_{j}^{(2J+1)}(t)$ can be constructed formally as follows. Assuming temporarily that the right-hand sides of equations (16) vanish for $11\leq j\leq J-$ 1, one gets $r_{j-1}r_{j+1}=r_{j}^{2}$, thus $r_{j}$ form a geometric sequence
@@
and therefore $r_{J-1}=r_{J}^{\frac{J-1}{J}}$. Substituting this into Eq. (16) for $j=J$ gives $\ddot{r}_{J}=-r_{J}^{-\frac{J+2}{J}},$ whose zero energy solution is $r_{J}(t)=(At)^{\frac{J}{J+1}}$. Combining this result with (20) yields (18). Having the leading order asymptotics, the solution can be systematically improved by adding subleading corrections. Let us illustrate how this goes for $J=2$, where (18) reads
@@ 4. PDE computations
In this section we solve Eq. (6) numerically using the hyperboloidal formulation of the initial value problem [13]. To this end we define a new time coordinate $s=t-\cosh{x}$. The hypersurfaces of constant $s$ are “hyperboloidal” which means that they are spacelike hypersurfaces that approach the future null infinities of the wormhole spacetime along the outgoing null cones for $x\rightarrow\pm\infty$. In terms of $s$ Eq. (6) becomes (for $k=2$)
@@
where $b_{\pm}(s)$ are the radiation coefficients. Multiplying Eq. (36) by $u_{s}$ and rearranging terms we obtain the conservation law
@@
Note that the potential part of the energy $\mathcal{E}$ is the same as $V(u)$ for the conserved energy $E$, however the kinetic part is different. Since the Bondi energy is monotonically decreasing and bounded from below—by 00 for $n=0$ and by $4$ for $n=1$ [due to Bogomolnyi’s inequality (8)]—there exists a limit
@@
In numerical studies we introduce the coordinate $y=\tanh\left(x/4\right)$ which compactifies the spatial domain to the interval $[-1,1]$ and transforms Eq. (36) to
$u_{ss}+\frac{2y(1+y^{2})}{1-y^{2}}\,u_{sy}+\frac{y^{4}+6y^{2}+1}{(1-y^{2})^{2}}u_{s}=\frac{1}{16}(1-y^{2})\partial_{y}\left[(1-y^{2})u_{y}\right]-2\sin(2u).$ (42)
@@
Acknowledgement. This work was supported by the National Science Centre grant No. 2017/26/A/ST2/00530 (P.B.), the ERC Starting Grant “INSOLIT” 101117126 (J.J.), the Austrian Science Fund (FWF) Project P 36455 and START-Project Y 963 (M.M.). M.M. acknowledges the hospitality and support of Laboratoire Analyse, Géométrie et Applications, Sorbonne Paris North University, during the research stay when part of this work was conducted.